OpenStudy (anonymous):

PbO2 oxidises MnO in presence of HNO3 the products are HMnO4 , Pb(nO3)2 and h2o calculate the number of moles of HNO3 and PbO2 consumed per mole of MnO?

sam (.sam.):

You just balance the eqation

OpenStudy (anonymous):

3 MnO + 5 PbO2 + 5 HNO3 = 3 HMnO4 + 5 Pb(NO3) + H2O use this equation

OpenStudy (anonymous):

the ans is 4 ,2

sam (.sam.):

\[\begin{array}{cc} 5 \text{Pb}\text{O}_2 \\\end{array}+2MnO+10HNO_3 \rightarrow 2MnO_4+\begin{array}{cc} 5 \text{Pb}\left(\text{N}\text{O}_3\right)_2 \\ \end{array}+4H_2O\]

OpenStudy (anonymous):

it is HMnO4

sam (.sam.):

I mistake, but the balancing is correct

sam (.sam.):

I've even checked with a software

OpenStudy (anonymous):

the answer is 4,2 plzzz check it

OpenStudy (anonymous):

calculate the number of moles of HNO3 and PbO2 consumed per mole of MnO?

sam (.sam.):

If per mole of MnO then you have 5 moles of HNO3 and 2.5Moles of PbO2

sam (.sam.):

you sure the reaction is correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@.Sam. plzz help

OpenStudy (anonymous):

calculate the number of moles of HNO3 and PbO2 consumed per mole of MnO? i think we have to do something here

sam (.sam.):

The balancing is \[\begin{array}{cc} 5 \text{Pb}\text{O}_2 \\\end{array}+2MnO+10HNO_3 \rightarrow 2HMnO_4+\begin{array}{cc} 5 \text{Pb}\left(\text{N}\text{O}_3\right)_2 \\ \end{array}+4H_2O\] If per mole of MnO then you have 5 moles of HNO3 and 2.5Moles of PbO2, should be correct

sam (.sam.):

you just divide 2 to all reactants and you get per mole of MnO

sam (.sam.):

then you get per mole of MnO then 5 moles of HNO3 and 2.5Moles of PbO2

OpenStudy (anonymous):

i think 4,2 is wrong answer.because when changing from MnO to HMnO4 Mn requires 5 electrons(since it changes from +2 to +7).each PbO2 will provide just 2 electrons.so 2 PbO2 will give only four electrons,which isn't sufficient.